You Do The Math: Calculating Energy Contribution of Carbohydrate, Fat, and Protein

Calculate the missing values in the table (Ques. A%, Ques. B% etc.) and then answer the bottom two questions (I and J).Calories CHO (g) FAT (g) PRO (g) 770 Menu/Food Menu 1/McDonalds Burger Meal Double Quarter Pounder with Cheese French Fries
Answer #1

Percentage of a nutrient in the food = (Total grams of nutrient * physiological value of the nutrient ) / Total value of Calories

Multiply the result by 100 to convert into percentage

Menu 1:

(A %) Percentage of CHO = (244*4) / 2190 = 0.445 = 44.5%

(B%) Percentage of FAT = (99*9) / 2190 = 0.406 = 40.6%

Menu 2:

(C%) Percentage of FAT = (21*9) / 620 = 0.304 = 30%

(D%) Percentage of PRO = (42*4) / 620 = 0.27 = 27%

Menu 3:

(E%) Percentage of FAT = (19*9) / 490 = 0.348 = 0.35 = 35%

(F%) Percentage of PRO = (15*4) / 490 = 0.12 = 12%

Menu 4:

(G%) Percentage of FAT = (47*9) / 1320 = 0.32 = 32%

(H%) Percentage of PRO = (34*4) / 1320 = 0.10 = 10%

Q I (Ans): Menu 3 is lowest in Carbohydrate total grams

63*4 = 252 grams

Q J ( Ans): Menu 2 is lowest in Carbohydrate Percentage

(68*4) / 620 = 0.438 = 0.44 = 44%